Support department cost allocation—sequential method

Crystal Scarves & Co. produces winter scarves. The scarves are produced in the Cutting and Sewing departments. The Maintenance and Security departments support these production departments, and allocate costs based on machine hours and square feet, respectively. Information about each department is provided in the following table:

Department Total Cost Number of Employees Machine Hours Square Feet
Maintenance Department $ 2,300 6    57    800
Security Department 3,000 4    0    600
Cutting Department 19,600 20    3,700    3,200
Sewing Department 20,800 18    5,550    4,000

Using the sequential method and allocating the support department with the highest costs first, allocate all support department costs to the production departments. Then compute the total cost of each production department.

Cutting
Department
Sewing
Department
Production departmentsʼ total costs $ $
Answer #1

Answer

Cutting Sewing
Producion department total costs $ 21,840 $ 23,860

Calculations:

Maintainance Security Cutting Sewing
Total Cost $                2,300 $   3,000 $ 19,600 $ 20,800
Security department $                   300 $ (3,000) $    1,200 $    1,500
Total $                2,600 $          -   $ 20,800 $ 22,300
Maintainance department $              (2,600) $          -   $    1,040 $    1,560
Total $                      -   $          -   $ 21,840 $ 23,860

Working Note:

Maintenance Security Cutting Sewing
Security department =3,000/8,000*800 -3,000 =3,000/8,000*3,200 =3,000/8,000*4,000
Maintenance department -2,600 0 =2,600/9,250*3,700 =2,600/9,250*5,550

In case of any doubt, please comment.